OpenStudy (anonymous):

Write this using radical notation. Assume that all variables represent positive real numbers. (5b)^(4/5)

jimthompson5910 (jim_thompson5910):

\[(5b)^{\frac{4}{5}}=\sqrt[5]{(5b)^4}=\sqrt[5]{625b^4}\]

OpenStudy (anonymous):

Thank you so much!